eBook Your client is 28 years old. She wants to begin saving for retirement, with the first payment to come one year from now. She can save $5,000 per year, and you advise her to invest it in the stock market, which you expect to provide an average return of 12% in the future. a. If she follows your advice, how much money will she have at 65? Do not round intermediate calculations. Round your answer to the nearest cent. b. How much will she have at 70? Do not round intermediate calculations. Round your answer to the nearest cent. c. She expects to live for 20 years if she retires at 65 and for 15 years if she retires at 70. If her investments continue to earn the same rate, how much will she be able to withdraw at the end of each year after retirement at each retirement age? Do not round intermediate calculations. Round your answers to the nearest cent. Annual withdrawals if she retires at 65: $ Annual withdrawals if she retires at 70: $
Answer #1

a.Number of years = 65 – 28 = 37 years

Total Amount of Money will be equal to future value of annuity

= 5,000*[{(1+0.12)37-1}/0.12]

= $27,17,993.45

b.At 70, Number of years = 42

Amount = 5,000*[{(1+0.12)42-1}/0.12]

= $48,21,797.39

c.Let she withdraws $x per year

Present Value of Amount withdrawn = $27,17,993.45

27,12,993.45 = x*PVAF(12%, 20 years)

27,12,993.45 = x*7.4694

X = $363,214.37

If retires at 70:

48,21,797.39 = x*PVAF(12%, 15 years)

48,21,797.39 = x*6.8109

X = $707,953.04